# Density Ratio Estimation-based Bayesian Optimization with Semi-Supervised Learning

### v3

@@ Abstract
Bayesian optimization has attracted huge attention from diverse research areas in science and engineering, since it is capable of efficiently finding a global optimum of an expensive-to-evaluate black-box function. In general, a probabilistic regression model is widely used as a surrogate function to model an explicit distribution over function evaluations given an input to estimate and a training dataset. Beyond the probabilistic regression-based methods, density ratio estimation-based Bayesian optimization has been suggested in order to estimate a density ratio of the groups relatively close and relatively far to a global optimum. Developing this line of research further, supervised classifiers are employed to estimate a class probability for the two groups instead of a density ratio. However, the supervised classifiers used in this strategy are prone to be overconfident for known knowledge on global solution candidates. Supposing that we have access to unlabeled points, e.g., predefined fixed-size pools, we propose density ratio estimation-based Bayesian optimization with semi-supervised learning to solve this challenge. Finally, we show the empirical results of our methods and several baseline methods in two distinct scenarios with unlabeled point sampling and a fixed-size pool and analyze the validity of our proposed methods in diverse experiments.
@@ 1 Introduction
Bayesian optimization (Brochu et al., 2010; Garnett, 2023) has attracted immense attention from various research areas such as hyperparameter optimization (Bergstra et al., 2011), battery lifetime optimization (Attia et al., 2020), chemical reaction optimization (Shields et al., 2021), and nanophotonic structure optimization (Kim et al., 2024), since it is capable of efficiently finding a global optimum of an expensive-to-evaluate black-box function. As studied in previous literature on Bayesian optimization (Snoek et al., 2012; Martinez-Cantin et al., 2018; Springenberg et al., 2016; Hutter et al., 2011), a probabilistic regression model, which can estimate a distribution of function evaluations over inputs, is widely used as a surrogate function; A Gaussian process (GP) (Rasmussen & Williams, 2006) is a predominant choice for the surrogate function. An analogy between probabilistic regression models in Bayesian optimization is that they rely on an explicit function over function evaluations $p(y\mid\mathbf{x},\mathcal{D})$ given an input to estimate, denoted as $\mathbf{x}$, and a training dataset $\mathcal{D}$.
@@
However, the supervised classifiers utilized in the DRE-based Bayesian optimization are prone to be overconfident for known knowledge on global solution candidates or a region that have been already exploited. In this paper, supposing that we have access to unlabeled points, we propose a novel DRE-based method with semi-supervised learning, which is named DRE-BO-SSL, to solve this over-exploitation problem. Although our direct competitors, i.e., BORE (Tiao et al., 2021) and LFBO (Song et al., 2022), show their strengths through theoretical and empirical analyses, our algorithm betters an ability to consider a wider region that satisfies $p(\mathbf{x}\mid y\leq y^{\dagger},\mathcal{D})\geq p(\mathbf{x}\mid y>y^{\dagger},\mathcal{D})$, than the competitors, as shown in Figure 1. By this intuitive example in Figure 1, we presume that DRE-BO-SSL appropriately balances exploration and exploitation rather than the existing methods. Compared to a supervised classifier, e.g., random forests (Breiman, 2001), gradient boosting (Friedman, 2001), and multi-layer perceptrons, our semi-supervised classifiers, i.e., label propagation (Zhu & Ghahramani, 2002) and label spreading (Zhou et al., 2003), are less confident in terms of the regions of global solution candidates by using unlabeled data points; see Figures 1 and 9 and Sections 1.1 and 3 for detailed examples and analyses.
To make use of semi-supervised classifiers, we take into account two distinct scenarios with unlabeled point sampling and with a predefined fixed-size pool. For the first scenario, we randomly sample unlabeled data points from the truncated multivariate normal distribution using a minimax tilting method (Botev, 2017), in order that it is possible that a cluster assumption (Seeger, 2000) is assumed. Finally, we demonstrate that our method shows superior performance compared to the exiting methods in diverse experiments including synthetic benchmarks, Tabular Benchmarks (Klein & Hutter, 2019), NATS-Bench (Dong et al., 2021), and 64D minimum multi-digit MNIST search. Note that Tabular Benchmarks, NATS-Bench, and 64D minimum multi-digit MNIST search have access to fixed-size pools. Furthermore, in order to validate our methods, we provide thorough analyses on the components of DRE-BO-SSL in Section 6 and the appendices.
To sum up, our contributions are itemized as follows:
We promise to release our implementation upon publication.
@@ 1.1 Over-exploitation Problem
As illustrated in Figures 1 and 9, the supervised classifiers used in DRE-based Bayesian optimization suffer from the over-exploitation problem. Interestingly, many deep learning models also share a similar problem, which is called an overconfidence problem (Guo et al., 2017; Müller et al., 2019), due to various reasons, but primarily due to overparameterization. It is noteworthy that the definition of the over-exploitation problem in DRE-based Bayesian optimization is different from the overconfidence problem in general classification. The definition in DRE-based Bayesian optimization does not imply that a single data point has a high probability for a particular class, but it implies that a few data points have high probabilities for a class of interest. More precisely, our definition indicates the problem of overconfidence over known knowledge on global solution candidates. For example, the definition in general classification includes a case that 50% of data points are biased to one class and the remainder is biased to another class. On the contrary, this definition does not include such a case and it only includes a case that a small number of data points or the small region of a search space, which is colored in yellow in Figures 1 and 9, are biased to some particular class, i.e., Class 1 in the DRE-based Bayesian optimization problems.
@@ 2 Background and Related Work
It is a principled and efficient approach to finding a global solution of a challenging objective, e.g., expensive-to-evaluate black-box functions (Brochu et al., 2010; Garnett, 2023). To focus on a probabilistic regression model as a surrogate function, we omit the details of Bayesian optimization here; see (Brochu et al., 2010; Garnett, 2023) for details. In Bayesian optimization, GP regression (Rasmussen & Williams, 2006) is widely used as a surrogate function (Srinivas et al., 2010; Snoek et al., 2012) because of its flexibility with minimum assumptions on model and smoothness. While GP is a probable choice, Bayesian optimization with diverse surrogates such as Student-$t$ process regression (Martinez-Cantin et al., 2018), Bayesian neural networks (Springenberg et al., 2016), and tree-based models (Hutter et al., 2011; Kim & Choi, 2022) has been proposed. An analogy between such models is that they model $p(y\mid\mathbf{x},\mathcal{D})$ explicitly, so that it can be used to define an acquisition function with the statistics of $p(y\mid\mathbf{x},\mathcal{D})$.
@@ 3 DRE-based Bayesian Optimization
$A(\mathbf{x}\!\mid\!\zeta,\mathcal{D}_{t})\!=\!\frac{p(\mathbf{x}\!\mid\!y\!\leq\!y^{\dagger},\mathcal{D}_{t})}{\zeta p(\mathbf{x}\!\mid\!y\!\leq\!y^{\dagger},\mathcal{D}_{t})\!+\!(1\!-\!\zeta)p(\mathbf{x}\!\mid\!y\!>\!y^{\dagger},\mathcal{D}_{t})},$ (1)
@@
By the Bayes’ theorem, (2) becomes $A(\mathbf{x}\mid\zeta,\mathcal{D}_{t})=\zeta^{-1}p(z=1\mid\mathbf{x})/(p(z=1\mid\mathbf{x})+p(z=0\mid\mathbf{x}))$. Therefore, a class probability over $\mathbf{x}$ for Class 1 is considered as an acquisition function; it is simply denoted as $A(\mathbf{x}\mid\zeta,\mathcal{D}_{t})=\zeta^{-1}\pi(\mathbf{x})$. Eventually, the class probability is estimated by various off-the-shelf classifiers such as random forests and multi-layer perceptrons.
@@
$A(\mathbf{x}\mid\zeta,\mathcal{D}_{t})=\operatornamewithlimits{\arg\max}_{S:\mathcal{X}\to\mathbb{R}}\mathbb{E}_{\mathcal{D}_{t}}[u(y;y^{\dagger})f^{\prime}(S(\mathbf{x}))-f^{*}(f^{\prime}(S(\mathbf{x})))],$ (3)
@@ 4.1 Label Propagation and Label Spreading
Suppose that each data point is defined on a $d$-dimensional compact space $\mathcal{X}\subset\mathbb{R}^{d}$. We consider $n_{l}$ labeled points $\mathbf{X}_{l}\in\mathbb{R}^{n_{l}\times d}$, their corresponding labels $\mathbf{C}_{l}\in\mathbb{R}^{n_{l}\times c}$, and $n_{u}$ unlabeled points $\mathbf{X}_{u}\in\mathbb{R}^{n_{u}\times d}$, where $c$ is the number of classes. $\mathbf{X}_{l}$ and $\mathbf{C}_{l}$ are query points that have already been evaluated and their class labels; we define $\mathbf{X}_{l}=[\mathbf{x}_{0},\ldots,\mathbf{x}_{n_{l}-1}]^{\top}$ for $\mathcal{D}_{t}=\{(\mathbf{x}_{i},y_{i})\}_{i=0}^{n_{l}-1}$. For the sake of brevity, the concatenated data points of $\mathbf{X}_{l}$ and $\mathbf{X}_{u}$ are defined as $\mathbf{X}=[\mathbf{X}_{l};\mathbf{X}_{u}]\in\mathbb{R}^{(n_{l}+n_{u})\times d}$. Note that in our problem $c=2$ because we address the problem with two classes.
As shown in Algorithm 2, we first initialize labels to propagate $\widehat{\mathbf{C}}\in\mathbb{R}^{(n_{l}+n_{u})\times 2}$; it is initialized as $\widehat{\mathbf{C}}=[\mathbf{c}_{0},\ldots,\mathbf{c}_{n_{l}-1},\mathbf{c}_{n_{l}},\ldots,\mathbf{c}_{n_{l}+n_{u}-1}]^{\top}$, where $\mathbf{c}_{0},\ldots,\mathbf{c}_{n_{l}-1}$ are one-hot representations $[\mathbf{C}_{l}]_{1:},\ldots,[\mathbf{C}_{l}]_{n_{l}:}$, and $\mathbf{c}_{n_{l}},\ldots,\mathbf{c}_{n_{l}+n_{u}-1}$ are zero vectors. Denote that $[\mathbf{C}]_{i:}$ is $i$th row of $\mathbf{C}$. Then, we compute a similarity between two data points $\mathbf{x}_{i},\mathbf{x}_{j}\in\mathcal{X}$, so that we compare all pairs in $\mathbf{X}$. For example, one of popular similarity functions, i.e., a radial basis function kernel, can be used:
@@
See Figure 4 for the results of learning $\beta$ and Section E for analysis on the effects of $\beta$. By (5), we compute a transition probability from $\mathbf{x}_{j}$ to $\mathbf{x}_{i}$ by $p_{ij}=w_{ij}/\sum_{k=1}^{n_{l}+n_{u}}w_{kj}$. Note that similarities $\mathbf{W}\in\mathbb{R}^{(n_{l}+n_{u})\times(n_{l}+n_{u})}$ and transition probabilities $\mathbf{P}\in\mathbb{R}^{(n_{l}+n_{u})\times(n_{l}+n_{u})}$ are defined, where $[\mathbf{W}]_{ij}=w_{ij}$ and $[\mathbf{P}]_{ij}=p_{ij}$. Moreover, by the definition of $p_{ij}$, $\mathbf{P}=\mathbf{D}^{-1}\mathbf{W}$, where $\mathbf{D}$ is a degree matrix whose diagonal entry is defined as $[\mathbf{D}]_{ii}=\sum_{j=1}^{n_{l}+n_{u}}[\mathbf{W}]_{ij}$.
@@
$\widehat{\mathbf{C}}_{t+1}\leftarrow\mathbf{P}^{\top}\widehat{\mathbf{C}}_{t}\quad\textrm{and}\quad[\widehat{\mathbf{C}}_{t+1}]_{i:}\leftarrow[\mathbf{C}_{l}]_{i:},$ (7)
@@ 5 Experiments
We test DRE-BO-SSL and baseline methods in the following optimization problems: synthetic benchmarks for a scenario with unlabeled point sampling, and synthetic benchmarks, Tabular Benchmarks (Klein & Hutter, 2019), NATS-Bench (Dong et al., 2021), and minimum multi-digit MNIST search for a scenario with a fixed-size pool. Note that Tabular Benchmarks, NATS-Bench, and minimum multi-digit MNIST search are defined with a fixed number of possible solution candidates, which implies that they are considered as combinatorial optimization problems. By following the previous work (Tiao et al., 2021; Song et al., 2022), we set a threshold ratio as $\zeta=0.33$ for all experiments; the effects of a threshold ratio $\zeta$ are analyzed in Sections 6 and H. To solve (4), we use L-BFGS-B (Byrd et al., 1995) with 1000 different initializations. All experiments are repeated 20 times with 20 fixed random seeds and the sample mean and the standard error of the sample mean are reported. Other missing details including the details of the competitors of our methods are presented in Section D.

### v4

@@ Abstract
Bayesian optimization has attracted huge attention from diverse research areas in science and engineering, since it is capable of efficiently finding a global optimum of an expensive-to-evaluate black-box function. In general, a probabilistic regression model is widely used as a surrogate function to model an explicit distribution over function evaluations given an input to estimate and a training dataset. Beyond the probabilistic regression-based methods, density ratio estimation-based Bayesian optimization has been suggested in order to estimate a density ratio of the groups relatively close and relatively far to a global optimum. Developing this line of research further, supervised classifiers are employed to estimate a class probability for the two groups instead of a density ratio. However, the supervised classifiers used in this strategy are prone to be overconfident for known knowledge on global solution candidates. Supposing that we have access to unlabeled points, e.g., predefined fixed-size pools, we propose density ratio estimation-based Bayesian optimization with semi-supervised learning to solve this challenge. Finally, we show the empirical results of our methods and several baseline methods in two distinct scenarios with unlabeled point sampling and a fixed-size pool, and analyze the validity of our methods in diverse experiments.
Bayesian Optimization, Density Ratio Estimation-based Bayesian Optimization, Bayesian Optimization with Semi-Supervised Learning
@@ 1 Introduction
Bayesian optimization (Brochu et al., 2010; Garnett, 2023) has attracted immense attention from various research areas such as hyperparameter optimization (Bergstra et al., 2011), battery lifetime optimization (Attia et al., 2020), chemical reaction optimization (Shields et al., 2021), nanophotonic structure optimization (Kim et al., 2024), and language model fine-tuning (Jang et al., 2024), since it is capable of efficiently finding a global optimum of an expensive-to-evaluate black-box function. As studied in previous literature on Bayesian optimization (Snoek et al., 2012; Martinez-Cantin et al., 2018; Springenberg et al., 2016; Hutter et al., 2011), a probabilistic regression model, which can estimate a distribution of function evaluations over inputs, is widely used as a surrogate function; A Gaussian process (Rasmussen & Williams, 2006) is a predominant choice for the surrogate function. An analogy between probabilistic regression models in Bayesian optimization is that they rely on an explicit function over function evaluations $p(y\mid\mathbf{x},\mathcal{D})$ given an input to estimate, denoted as $\mathbf{x}$, and a training dataset $\mathcal{D}$.
@@
However, the supervised classifiers utilized in the DRE-based Bayesian optimization are prone to be overconfident for known knowledge on global solution candidates or a region that have been already exploited. In this paper, supposing that we have access to unlabeled points, we propose a novel DRE-based method with semi-supervised learning, which is named DRE-BO-SSL, to solve this over-exploitation problem. Although our direct competitors, i.e., BORE (Tiao et al., 2021) and LFBO (Song et al., 2022), show their strengths through theoretical and empirical analyses, our algorithm betters an ability to consider a wider region that satisfies $p(\mathbf{x}\mid y\leq y^{\dagger},\mathcal{D})\geq p(\mathbf{x}\mid y>y^{\dagger},\mathcal{D})$, than the competitors, as shown in Figure 1. By this intuitive example in Figure 1, we presume that DRE-BO-SSL appropriately balances exploration and exploitation rather than the existing methods. Compared to a supervised classifier, e.g., random forests (Breiman, 2001), gradient boosting (Friedman, 2001), and multi-layer perceptrons, our semi-supervised classifiers, i.e., label propagation (Zhu & Ghahramani, 2002) and label spreading (Zhou et al., 2003), are less confident in terms of the regions of global solution candidates using unlabeled data points; see Figures 1 and 8 and Sections 1.1 and 3 for detailed examples and analyses.
To make use of semi-supervised classifiers, we take into account two distinct scenarios with unlabeled point sampling and with a predefined fixed-size pool. For the first scenario, we randomly sample unlabeled data points from the truncated multivariate normal distribution using a minimax tilting method (Botev, 2017), in order that it is possible that a cluster assumption (Seeger, 2000) is assumed. Finally, we demonstrate that our method shows superior performance compared to the exiting methods in diverse experiments including synthetic benchmarks, Tabular Benchmarks (Klein & Hutter, 2019), NATS-Bench (Dong et al., 2021), and 64D minimum multi-digit MNIST search. Note that Tabular Benchmarks, NATS-Bench, and 64D minimum multi-digit MNIST search have access to fixed-size pools. To validate our methods, we provide thorough analyses on the components of DRE-BO-SSL in Section 6 and the appendices.
To sum up, our contributions are itemized as follows:
@@ 1.1 Over-Exploitation Problem
As illustrated in Figures 1 and 8, the supervised classifiers used in DRE-based Bayesian optimization suffer from the over-exploitation problem. Interestingly, many deep learning models also share a similar problem, which is called an overconfidence problem (Guo et al., 2017; Müller et al., 2019), due to various reasons, but primarily due to overparameterization. It is noteworthy that the definition of the over-exploitation problem in DRE-based Bayesian optimization is different from the overconfidence problem in general classification. The definition in DRE-based Bayesian optimization does not imply that a single data point has a high probability for a particular class, but it implies that a few data points have high probabilities for a class of interest. More precisely, our definition indicates the problem of overconfidence over known knowledge on global solution candidates. For example, the definition in general classification includes a case that 50% of data points are biased to one class and the remainder is biased to another class. On the contrary, this definition does not include such a case and it only includes a case that a small number of data points or the small region of a search space, which is colored in yellow in Figures 1 and 8, are biased to some particular class, i.e., Class 1 in DRE-based Bayesian optimization.
@@ 2 Background and Related Work
It is a principled and efficient approach to finding a global solution of a challenging objective, e.g., expensive-to-evaluate black-box functions (Brochu et al., 2010; Garnett, 2023). To focus on a probabilistic regression model as a surrogate function, we omit the details of Bayesian optimization here; see (Brochu et al., 2010; Garnett, 2023) for details. In Bayesian optimization, Gaussian process regression (Rasmussen & Williams, 2006) is widely used as a surrogate function (Srinivas et al., 2010; Snoek et al., 2012) because of its flexibility with minimum assumptions on model and smoothness. While a Gaussian process is a probable choice, Bayesian optimization with diverse surrogates such as Student-$t$ process regression (Martinez-Cantin et al., 2018), Bayesian neural networks (Springenberg et al., 2016), and tree-based models (Hutter et al., 2011; Kim & Choi, 2022) has been proposed. An analogy between such models is that they model $p(y\mid\mathbf{x},\mathcal{D})$ explicitly, so that it can be used to define an acquisition function with the statistics of $p(y\mid\mathbf{x},\mathcal{D})$.
@@ 3 DRE-based Bayesian Optimization
$A(\mathbf{x}\mid\zeta,\mathcal{D}_{t})$
$=\frac{p(\mathbf{x}\mid y\leq y^{\dagger},\mathcal{D}_{t})}{\zeta p(\mathbf{x}\mid y\leq y^{\dagger},\mathcal{D}_{t})+(1-\zeta)p(\mathbf{x}\mid y>y^{\dagger},\mathcal{D}_{t})},$ (1)
@@
By the Bayes’ theorem, (2) becomes the following:
$A(\mathbf{x}\mid\zeta,\mathcal{D}_{t})=\zeta^{-1}\frac{p(z=1\mid\mathbf{x})}{p(z=1\mid\mathbf{x})+p(z=0\mid\mathbf{x})}.$ (3)
Therefore, a class probability over $\mathbf{x}$ for Class 1 is considered as an acquisition function; it is simply denoted as the following:
$A(\mathbf{x}\mid\zeta,\mathcal{D}_{t})=\zeta^{-1}\pi(\mathbf{x}).$ (4)
Eventually, the class probability is estimated by various off-the-shelf classifiers such as random forests and multi-layer perceptrons.
@@
$A(\mathbf{x}\mid\zeta,\mathcal{D}_{t})$
$=\operatornamewithlimits{\arg\max}_{S:\mathcal{X}\to\mathbb{R}}\mathbb{E}_{\mathcal{D}_{t}}[u(y;y^{\dagger})f^{\prime}(S(\mathbf{x}))-f^{*}(f^{\prime}(S(\mathbf{x})))],$ (5)
@@ 4.1 Label Propagation and Label Spreading
Suppose that each data point is defined on a $d$-dimensional compact space $\mathcal{X}\subset\mathbb{R}^{d}$. We consider $n_{l}$ labeled points $\mathbf{X}_{l}\in\mathbb{R}^{n_{l}\times d}$, their corresponding labels $\mathbf{C}_{l}\in\mathbb{R}^{n_{l}\times c}$, and $n_{u}$ unlabeled points $\mathbf{X}_{u}\in\mathbb{R}^{n_{u}\times d}$, where $c$ is the number of classes. $\mathbf{X}_{l}$ and $\mathbf{C}_{l}$ are query points that have already been evaluated and their class labels; we define $\mathbf{X}_{l}=[\mathbf{x}_{0},\ldots,\mathbf{x}_{n_{l}-1}]^{\top}$ for $\mathcal{D}_{t}=\{(\mathbf{x}_{i},y_{i})\}_{i=0}^{n_{l}-1}$. For the sake of brevity, the concatenated data points of $\mathbf{X}_{l}$ and $\mathbf{X}_{u}$ are defined as $\mathbf{X}=[\mathbf{X}_{l};\mathbf{X}_{u}]\in\mathbb{R}^{(n_{l}+n_{u})\times d}$. Note that in our problem $c=2$, because we address the problem with two classes.
As shown in Algorithm 2, we first initialize labels to propagate $\widehat{\mathbf{C}}\in\mathbb{R}^{(n_{l}+n_{u})\times 2}$; it is initialized as the following:
$\widehat{\mathbf{C}}=[\mathbf{c}_{0},\mathbf{c}_{1},\ldots,\mathbf{c}_{n_{l}-1},\mathbf{c}_{n_{l}},\mathbf{c}_{n_{l}+1},\ldots,\mathbf{c}_{n_{l}+n_{u}-1}]^{\top},$ (7)
where $\mathbf{c}_{0},\mathbf{c}_{1},\ldots,\mathbf{c}_{n_{l}-1}$ are one-hot representations $[\mathbf{C}_{l}]_{1:},[\mathbf{C}_{l}]_{2:},\ldots,[\mathbf{C}_{l}]_{n_{l}:}$, and $\mathbf{c}_{n_{l}},\mathbf{c}_{n_{l}+1},\ldots,\mathbf{c}_{n_{l}+n_{u}-1}$ are zero vectors. Denote that $[\mathbf{C}]_{i:}$ is $i$th row of $\mathbf{C}$. Then, we compute a similarity between two data points $\mathbf{x}_{i},\mathbf{x}_{j}\in\mathcal{X}$, so that we compare all pairs in $\mathbf{X}$. For example, one of popular similarity functions, i.e., a radial basis function kernel, can be used:
@@
See Figure 4 for the results of learning $\beta$ and Section E for analysis on the effects of $\beta$. By (8), we compute a transition probability from $\mathbf{x}_{j}$ to $\mathbf{x}_{i}$ by $p_{ij}=w_{ij}/\sum_{k=1}^{n_{l}+n_{u}}w_{kj}$. Note that similarities $\mathbf{W}\in\mathbb{R}^{(n_{l}+n_{u})\times(n_{l}+n_{u})}$ and transition probabilities $\mathbf{P}\in\mathbb{R}^{(n_{l}+n_{u})\times(n_{l}+n_{u})}$ are defined, where $[\mathbf{W}]_{ij}=w_{ij}$ and $[\mathbf{P}]_{ij}=p_{ij}$. Moreover, by the definition of $p_{ij}$, $\mathbf{P}=\mathbf{D}^{-1}\mathbf{W}$, where $\mathbf{D}$ is a degree matrix whose diagonal entry is defined as the following:
$[\mathbf{D}]_{ii}=\sum_{j=1}^{n_{l}+n_{u}}[\mathbf{W}]_{ij}.$ (10)
@@
$\widehat{\mathbf{C}}_{t+1}$ $\leftarrow\mathbf{P}^{\top}\widehat{\mathbf{C}}_{t},$ (11)
$[\widehat{\mathbf{C}}_{t+1}]_{i:}$ $\leftarrow[\mathbf{C}_{l}]_{i:},$ (12)
@@ 5 Experiments
We compare baseline methods with DRE-BO-SSL in the following optimization problems: synthetic benchmarks for a scenario with unlabeled point sampling, and synthetic benchmarks, Tabular Benchmarks (Klein & Hutter, 2019), NATS-Bench (Dong et al., 2021), and minimum multi-digit MNIST search for a scenario with a fixed-size pool. Note that Tabular Benchmarks, NATS-Bench, and minimum multi-digit MNIST search are defined with a fixed number of possible solution candidates, which implies that they are considered as combinatorial optimization problems. By following the previous work (Tiao et al., 2021; Song et al., 2022), we set a threshold ratio as $\zeta=0.33$ for all experiments; the effects of a threshold ratio $\zeta$ are analyzed in Sections 6 and H. To solve (6), we use L-BFGS-B (Byrd et al., 1995) with 1,000 different initializations. All experiments are repeated 20 times with 20 fixed random seeds, where 5 initial points are given to each experiment. The sample mean and the standard error of the sample mean are reported. Other missing details including the details of the competitors of our methods are presented in Section D.
  • Gaussian process, EI and UCB: It is a Bayesian optimization strategy, which is defined with Gaussian process regression with the Matérn 5/2 kernel (Rasmussen & Williams, 2006), where expected improvement (Jones et al., 1998) or GP-UCB (Srinivas et al., 2010) is used as an acquisition function;
  • Random forest, BORE and LFBO: These are BORE and LFBO that employ random forests (Breiman, 2001) with 1,000 decision trees, where minimum samples to split are set to 2 for these baselines;
  • Gradient boosting, BORE and LFBO: These methods are BORE and LFBO with gradient boosting classifiers (Friedman, 2001) with 100 decision trees, where a learning rate for the classifier is set to 0.3;
  • XGBoost, BORE and LFBO: Similar to gradient boosting, BORE and LFBO with XGBoost (Chen & Guestrin, 2016) have 100 decision trees as base learners with a learning rate of 0.3;
  • MLP, BORE and LFBO: These methods are built with two-layer fully-connected networks; the detail of the multi-layer perceptron is described as follows.
The architecture of the multi-layer perceptron is set as the following: (i) first layer: fully-connected, input dimensionality $d$, output dimensionality 32, ReLU; (ii) second layer: fully-connected, input dimensionality 32, output dimensionality 1, Logistic, where $d$ is the dimensionality of the problem we solve.
Note that most configurations for the baselines follow the configurations described in the work (Song et al., 2022).
(a) Naval
(b) Protein
(c) Parkinson’s
(d) Slice